OpenStudy (anonymous):

The size of an exponentially growing bacteria colony doubles in 2 hours. How long will it take for the number of bacteria to triple? Give your answer in exact form and decimal form. Exact form

OpenStudy (anonymous):

Exponential growth and the double growth indicates the population size is modeled by \[y=a(2^{rx})\] with \(a\) being the initial population, \(r\) being the relative growth factor, and \(x\) is time. The population doubles in 2 hours, so when \(x=2\), you have \[2a=a2^{2r}\\ 2=2^{2r}\\ 1=2^{2r-1}\\ r=\frac{1}{2}\] So you have to find \(x\) such that \(y=3a\): \[3a=a\left(2^{x/2}\right)\] Sounds right, at any rate. I might be wrong.

OpenStudy (wolf1728):

If it doubles every 2 hours then formula is 2^((t-2)/2) you want to solve when 2^((t-2)/2) = 3

OpenStudy (anonymous):

@wolf1728 So i solve 2^((t-2)/2)=3 ?

OpenStudy (anonymous):

for t?

OpenStudy (anonymous):

 Is this right?